OpenStudy (anonymous):

calc 3 help evaluate the line integral (y-x)dx+xydy along the line segment from (3,4) to (2,1)

OpenStudy (anonymous):

Well find the equation of the line from (3,4) to (2,1). We do not need an extra parameter in this case, we may simply write it in terms of x: y=3x-5. And so: dy=3dx. Now we are integrating along the above line from x=2 to x=3. Plug all of this in: \[\int\limits_{2}^{3}[(3x-5-x)+x(3x-5)(3)]dx=\int\limits_{2}^{3}(9x^2-13x-5)dx=\left [3x^3-\frac{13}{2}x^2-5x\right ]_2^3\]Which is equal to 39.